# TerraMind: Large-Scale Generative Multimodality for Earth Observation

### v1

@@ 5.6 Thinking in modalities
We employ the TiM tuning approach on water body mapping and during finetuning use radar or optical data to generate additional artificial LULC data which we leverage along the input. Our results in Table 7 indicate a superior performance of the TiM finetuning compared to leveraging uni-modal data by up to 2pp mIoU. This finding points us in the direction of TerraMind being able to generate data that improve finetuning to downstream applications.
@@ 6 Conclusion
7 Acknowledgements
This work is part of the Fostering Advancements in Foundation Models via Unsupervised and Self-supervised Learning for Downstream Tasks in Earth Observation (FAST-EO) project, which is funded by the European Space Agency (ESA) $\Phi$-lab under contract No. 4000143501/23/I-DT. The authors gratefully acknowledge the Gauss Centre for Supercomputing e.V. (www.gauss-centre.eu) for funding this project by providing computing time on the GCS Supercomputer JUWELS [3] at Jülich Supercomputing Centre (JSC).
@@ 8 TerraMesh Dataset
Figure 8: Visualization of the spatial-temporal alignment across modalities in TerraMesh. S-2 L2A uses IRRG pseudo-coloring and S-1 RTC is visualized as VH-VV-VH. To account for their value ranges, VV is scaled from -30 – 5 dB while VH is scaled from -40 – 0 dB. Copernicus DEM is scaled based on the image value range.
@@ 10 Tokenizer performance and general learnings
In Figures 10 to 12, we provide an overview on the spatial distributions of the S-1 GRD, S-2 L2A, and DEM tokenizer on the validation data of the SSL4EO-S12 subset which is focused on urban areas and therefore relevant for many downstream applications. Overall, we observe low MSE errors and particularly low deviation across geographic regions. For optical S-2 data, we observe minor difficulties in reconstructing images from Northern Asia, which we manually investigated. Overall, the vast majority of those samples are depicting snowy/icy conditions that have very high reflectance values of up to 12,000 compared to a normal range of [0, 255] in RGB data. On those long tail distribution samples, the S-2 tokenizer naturally has more difficulties.
@@
We give a more extensive background on radar data in the following for interested readers and non-EO experts. Reconstructing realistic and accurate synthetic aperture radar (SAR) S-1 VV and VH data is challenging due to factors inherent in the specific characteristics of SAR and the S-1 mission. SAR data is affected by complex interactions between the radar signal and Earth’s surface. SAR is based on radar backscatter, which is influenced by surface roughness and moisture content. The interaction of radar waves with different surfaces, including vegetation structure and urban environments, can produce complex backscatter patterns. The two polarizations, VV and VH, capture different scattering mechanisms: VV is sensitive to surface roughness and vegetation, while VH captures cross-polarized interactions that are influenced by surface and volumetric features [14, 56, 35]. In addition, SAR inherently contains speckle noise, which obscures fine details, making it difficult to extract accurate information. To evaluate the SAR data tokenizers of TerraMind, we employ various evaluation metrics to assess quality and accuracy. We compute the MSE for quantifying pixel-level differences, the SSIM to compare image structural content, and the PSNR [72, 67, 1].
We present the results in Table 9. Overall, we observe high to very high values in both SSIM (generally above 90% SSIM) and PSNR (generally above 55dB). These results validate the high quality of the encoding and decoding of the S-1 data using our tokenizers. Furthermore, we observe that the reconstructions are slightly better for the VV component of S-1 compared to the VH component. For reasons of comparison, we additionally provide the results for the S-2 tokenizer, which exhibits slightly lower SSIM score and strong PSNR scores. We additionally perform qualitative evaluations by visual inspection to identify artifacts, noise and other inconsistencies that may not be captured by numerical evaluation metrics on their own.
In addition to MSE, SSIM and PSNR in Table 9, for TerraMindv1-B, for the tokenizers GRD VV and VH, the correlation between the estimated and ground truth intensity values is $0.625$ and $0.701$, respectively, while for TerraMindv1-B-single, the correlations are $0.576$ and $0.670$. This improvement is likely related to the fact that the v1-B tokenizers have been trained on more (i.e. in number) and also much more diverse data. Moreover, the correlation between MSE and SSIM for all the test data samples is $-0.613$, which is interesting as SSIM is a different metric from MSE, while the correlation between MSE and PSNR is $-0.928$, i.e. closer to $-1$. These correlations are for S-1 GRD VV, while for VH, the correlations are $-0.416$ and $-0.929$.
TerraMindv1-B TerraMindv1-B-single
Tokenizer MSE SSIM PSNR MSE SSIM PSNR
S-1 GRD VV 0.118 0.961 58.1 0.154 0.961 57.8
S-1 GRD VH 0.172 0.940 56.2 0.207 0.939 56.0
S-2 L2A 0.037 0.926 66.3 0.032 0.933 66.8
@@ 12.1 Few-shot experiments
CLIP-ViT-B/16 S-2 RGB 52.39 74.39 39.71 62.80 28.13 37.51 22.21 30.92
CLIP-ViT-B/16 NAIP 31.73 42.84 24.86 36.41
@@ 12.2 Finetuning comparisons with baseline models
Table 11: Performance evaluation of TerraMind across seven benchmark datasets using the PANGAEA evaluation protocol. Higher mIoU values ($\uparrow$) and lower rank values ($\downarrow$) are reported. The best model per column is highlighted in bold and the second best is underscored. We indicate unimodal datasets with , demonstrating that TerraMind outperforms other benchmarks in both unimodal and multimodal settings.
Since the first approaches to foundation models for Earth observations, experts in the field discuss on the usability of such models compared to task-specific models that are trained for each application individually. Recent benchmark results suggested that task-specific models, like U-Nets, often outperform finetuned GFMs [49]. We therefore additionally investigate how TerraMind compares with task-specific U-Nets and ViT models following the PANGAEA evaluation protocol. As adviced by the authors of PANGAEA, we again report results on nine of the eleven datasets as we could not reproduce the performance on the remaining two datasets. The task-specific models are trained from scratch for each individual task, while all GFMs including TerraMind are finetuned with a frozen encoder and an UperNet head. For reasons of consistency, we report TerraMind results based on full finetuning instead of TiM tuning. Yet, for several tasks, we expect slightly higher results when applying TiM tuning for TerraMind. Overall, our results demonstrate that TerraMindv1-B outperforms task-specific UNet and ViT models across the PANGAEA benchmark in both unimodal and multimodal settings by 1pp avg. mIoU and 4pp avg. mIoU respectively. In multimodal settings, the improvement peaks to 4.5pp improvement of TerraMindv1-B over task-specific U-Nets. To the best of our knowledge, this is the first time a GFM model outperforms task-specific models on a global benchmark.

### v2

@@ 5.6 Thinking in modalities
We employ the TiM tuning approach on water body mapping and during finetuning use radar or optical data to generate additional artificial LULC data which we leverage along the input. Our results in Table 8 indicate a superior performance of the TiM finetuning compared to leveraging uni-modal data by up to 2pp mIoU. We also note that TiM tuning strongly benefits from a trainable encoder. This finding points us in the direction of TerraMind being able to generate data that improve finetuning to downstream applications.
@@ 7 TerraMesh Dataset
Figure 8: Visualization of the spatial-temporal alignment across modalities in TerraMesh. S-2 L2A uses IRRG pseudo-coloring and S-1 RTC is visualized in db scale as VH-VV-VV/VH. Copernicus DEM is scaled based on the image value range.
@@ 8.1 Tokenizer models
The tokenizer models are pretrained using a Vision Transformer (ViT) encoder and a patched UNet decoder, with input images ranging from 224x224 to 256x256 in size. The model was trained with patch sizes of 16x16 for the ViT encoder and 4x4 for the UNet decoder. A tanh MLP was used before the quantizer, as outlined in the ViT-VQGAN paper, to enhance tokenization quality.
The model utilized a Finite-Scalar Quantization (FSQ) approach with a codebook size of 8-8-8-6-5, aiming to learn consistent and abstract representations across image patches. The latent dimension was set to 5. We leverage the normalization of codebook entries to the unit sphere during training. This concept is borrowed from the ViT-VQGAN approach, which applies a specific form of normalization to improve the quality and efficiency of learned representations. Additionally, an EMA-based quantizer was used with a decay rate of 0.99 to track and improve quantization over time.
During diffusion-based pretraining, the model was trained for 1000 timesteps using a linear beta schedule, with MSE loss as the objective. The training leveraged half-precision (fp16) and used an AdamW optimizer with specific learning rate scheduling and warmup strategies. The model also incorporated model EMA for stable training and set a batch size of 1 per GPU with various regularization techniques like grad clipping and random horizontal flips.
@@ 9 Tokenizer performance and general learnings
9.1 FSQ vs. VQ
Generally, our pretraining experiments comparing FSQ with vector quantization suggest that both approaches can achieve the same level of performance, yet reaching optimal levels of performance with VQ is regarded to be more challenging than using FSQ. We visualize this through (a) the reconstruction loss and (b) the gradient norms of the tokenizer pretraining on S-2 L2A data in Figures 10 and 11, respectively. Overall, we observe that both approaches reach the same level of convergence, however FSQ requires less tuning and is generally more stable than VQ. This especially also applies for the grad norms.
Figure 10: Pretraining reconstruction losses of S-2 L2A modality comparing finite-scalar quantization (FSQ) and vector quantization (VQ) approaches. Overall, both approaches reach the same level of performance. The FSQ approach converges smoother than VQ, while requiring less tuning.
Figure 11: Gradient norms for pretraining of S-2 L2A tokenizers comparing finite-scalar quantization (FSQ) and vector quantization (VQ) approaches. The FSQ approach converges smoother than VQ, while requiring less tuning.
@@
In Figures 12 to 14, we provide an overview on the spatial distributions of the S-1 GRD, S-2 L2A, and DEM tokenizer on the validation data of the SSL4EO-S12 subset which is focused on urban areas and therefore relevant for many downstream applications. Overall, we observe low MSE errors and particularly low deviation across geographic regions. For optical S-2 data, we observe minor difficulties in reconstructing images from Northern Asia, which we manually investigated. Overall, the vast majority of those samples are depicting snowy/icy conditions that have very high reflectance values of up to 12,000 compared to a normal range of [0, 255] in RGB data. On those long tail distribution samples, the S-2 tokenizer naturally has more difficulties.
@@
We give a more extensive background on radar data in the following for interested readers and non-EO experts. Reconstructing realistic and accurate synthetic aperture radar (SAR) S-1 VV and VH data is challenging due to factors inherent in the specific characteristics of SAR and the S-1 mission. SAR data is affected by complex interactions between the radar signal and Earth’s surface. SAR is based on radar backscatter, which is influenced by surface roughness and moisture content. The interaction of radar waves with different surfaces, including vegetation structure and urban environments, can produce complex backscatter patterns. The two polarizations, VV and VH, capture different scattering mechanisms: VV is sensitive to surface roughness and vegetation, while VH captures cross-polarized interactions that are influenced by surface and volumetric features [14, 56, 35]. In addition, SAR inherently contains speckle noise, which obscures fine details, making it difficult to extract accurate information. To evaluate the SAR data tokenizers of TerraMind, we employ various evaluation metrics to assess quality and accuracy. We compute the MAE and RMSE for quantifying pixel-level differences, the SSIM to compare image structural content, and the PSNR [73, 67, 1].
Table 10 presents the quantitative evaluation of the TerraMind tokenizer reconstructions across multiple modalities. The results show a reasonable reconstruction performance for optical data, indicating both structural and perceptual fidelity. For radar modalities, S-1 GRD and S-1 RTC achieve comparable PSNR values, though SSIM scores are lower, suggesting that while the reconstructions are visually plausible, they exhibit moderate structural deviations. In addition to these quantitative metrics, we also conducted qualitative assessments through visual inspection to identify artifacts and inconsistencies not captured by numerical scores alone.
Modality MAE RMSE SSIM PSNR
S-1 GRD 2.403 3.220 0.565 30.291
S-1 RTC 2.216 2.888 0.466 30.389
S-2 L2A 0.055 0.134 0.851 27.439
DEM 170.7 737.2 0.974 20.712
NDVI 0.091 0.168 0.647 21.517
@@ 10 Additional experiments
10.1 Geolocation prediction
To better understand how TerraMind assigns geolocations, we further employ a Monte-Carlo sampling on the latitude-longitude grid for an optical tile from the validation data in Figure 15. We observe that while TerraMind is not predicting the correct geolocation ($\bullet$), there is a very high likelihood that the predicted geolocation is one of the adjacent grid points that have been seen during pretraining ($\bullet$). This result suggests that even for data from unseen geolocations, TerraMind remembers similar samples from the pretraining data ($\bullet$) and returns the geolocation of the samples with high similarity. This capability paired with the global pretraining of TerraMind suggests that geo-localization of data from unseen locations is possible but determined by the similarity to images from adjacent locations.
Figure 15: Distribution of predicted geo-locations for an optical S-2 L2A sample from the validation set. $\bullet$ is the correct location, $\bullet$ are Monte-Carlo sampled locations from TerraMind, $\bullet$ represents the distribution of training locations. TerraMind’s geo-localization seems to be based on similar optical samples in the training dataset for which TerraMind then outputs the geolocation.
We further extend the analysis of Figure 7 by additionally prompting the model for likely locations of urban areas. Overall, we observe that the model correctly identifies many densly populated areas across the globe. We also note overpredictions in, for example, North Africa and middle-east. This observation suggests that the model might confuse bare land and urban areas in these regions.
Figure 16: Prediction distribution of the land use class “urban” with a sampling temperature of $T=1.0$. TerraMind has a reasonable internal representation of the geolocation of specific contexts, like land use classes.
@@ 10.2 Few-shot experiments
CLIP-ViT-B/16 S-2 RGB 57.00 70.72 43.92 58.30 29.15 37.44 23.13 30.53
CLIP-ViT-B/16 NAIP 32.01 42.35 25.66 35.81
@@ 10.3 Finetuning comparisons with baseline models
Since the first approaches to foundation models for Earth observations, experts in the field discuss on the usability of such models compared to task-specific models that are trained for each application individually. Recent benchmark results suggested that task-specific models, like U-Nets, often outperform finetuned GFMs [49]. We therefore additionally investigate how TerraMind compares with task-specific U-Nets and ViT models following the PANGAEA evaluation protocol in Table 6. As adviced by the authors of PANGAEA, we again report results on nine of the eleven datasets as we could not reproduce the performance on the remaining two datasets. The task-specific models are trained from scratch for each individual task, while all GFMs including TerraMind are finetuned with a frozen encoder and an UperNet head. For reasons of consistency, we report TerraMind results based on full finetuning instead of TiM tuning. Yet, for several tasks, we expect slightly higher results when applying TiM tuning for TerraMind. Overall, our results demonstrate that TerraMindv1-B outperforms task-specific UNet and ViT models across the PANGAEA benchmark in both unimodal and multimodal settings by 1pp avg. mIoU and 4pp avg. mIoU respectively. In multimodal settings, the improvement peaks to 4.5pp improvement of TerraMindv1-B over task-specific U-Nets. To the best of our knowledge, this is the first time a GFM model outperforms task-specific models on a global benchmark.
@@ 10.5 TiM tuning for crop mapping
11 Any-to-any generation
Next to the evaluation of raw, pixel-level input in Table 3, we further evaluate the generation quality using tokenized input in Table 14. Interestingly, we observe only minor reduction in performance compared to pixel-level input even though the tokenized representations are compressed significantly (up to 3000x for S-2 L2A). Overall, our results suggest that leveraging tokenized inputs can be a reasonable alternative to leveraging pixel-level data for the generation of artificial data with TerraMind.
Modalities MAE RMSE SSIM PSNR
Tokenized S-2 L2A $\rightarrow$ S-1 GRD 3.3180 4.3309 0.5131 27.715
Tokenized S-2 L2A $\rightarrow$ S-1 RTC 3.0544 3.9178 0.4131 27.739
Tokenized S-2 L2A $\rightarrow$ DEM 572.5 1040.6 0.5728 17.718
Tokenized S-1 GRD $\rightarrow$ S-2 L2A 0.0820 0.1238 0.7182 25.630
Tokenized S-1 GRD $\rightarrow$ NDVI 0.1949 0.2425 0.4124 18.324
Tokenized S-1 GRD $\rightarrow$ DEM 327.4 550.3 0.7271 16.008
Tokenized S-1 RTC $\rightarrow$ S-2 L2A 0.1195 0.1935 0.6638 24.266
Tokenized S-1 RTC $\rightarrow$ NDVI 0.1895 0.2348 0.4500 18.606
Tokenized S-1 RTC $\rightarrow$ DEM 457.9 851.6 0.7095 19.457
Table 14: Performance of TerraMind on tokenized inputs using 10 diffusion steps. Metrics include MAE $\downarrow$, RMSE $\downarrow$, PSNR $\uparrow$, and SSIM $\uparrow$.
11.1 Large-scale generations
In Figures 19 and 21, we provide additional qualitative results for large-tile generations at the example of Singapore. Specifically, we leverage a 35.5km $\times$ 69.5km optical S-2 L2A tile as input and iteratively generate overlapping 224x224 pixel generations for S-1 RTC, S-1 GRD, NDVI, and LULC. In the overlapping areas, we apply the mean of all generations in order to enhance the spatial conciseness of the generations. TerraMind consistently removes the clouds in the S-1 generations. It makes assumptions for hidden areas, which are look accurate for large features like water bodies or the shore line. Other features like airports or ships are also clearly visible in the S-1 and NDVI generations.
(a) Input: S-2 L2A data from Singapore captured January 9th, 2025.
(b) Generation: TerraMind output for S-1 composition
Figure 19: Large-tile generations of TerraMind for Singapore (1/1)
(a) Generation: TerraMind output for LULC
Figure 20: Large-tile generations of TerraMind for Singapore (2/2)
(a) Input: S-2 L2A data from Santiago de Compostela.
(b) Generation: TerraMind output for S-1 GRD composition
Figure 21: Large-tile generations of TerraMind for Santiago de Compostela (1/3)
(a) TerraMind generation for S-1 RTC composition
(b) Generation: TerraMind output for vegetation
Figure 22: Large-tile generations of TerraMind for Santiago de Compostela (2/3)
(a) Generation: TerraMind output for digital elevation
Figure 23: Large-tile generations of TerraMind for Santiago de Compostela (3/3)